# A Dark Matter Fermionic Quantum Fluid from Standard Model Dynamics

### v2

@@ V Discussion and Conclusion
In this work we have presented an analysis regarding whether dark matter states are viable within the structure of the Standard Model. Although it is not possible, due to precision data, to construct dark states using only Standard Model fields, it is possible to do so within the gauge structure of the Standard Model, by augmenting the matter sector with well-motivated additional fermionic fields, namely the sterile neutrino and the vector-like quark. Each scenario has its unique aspects, but much of the physics common to both scenarios arises from the well-known physics of superconductivity.
@@
The necessary chemical potential in the neutrino case is adjusted to account for the fact that the neutrino mass is bounded from above and that the interaction is mediated by the Higgs which is relatively heavy for such an interaction. Due to the constraints on the neutrino mass and the Higgs mass, the possible values for the DM particle are fixed primarily by the chemical potential and it is thus possible in this case to generate DM masses in the ultralight range. In the vector-like quark case, because of the higher possible VLQ mass values and the massless (or nearly so) nature of the mediating particle, the scenario is not as strongly tuned and can naturally accommodate a wide range of DM masses from the ultralight to WIMP-scale DM.
@@
An important aspect left for future work is the phenomenology of this model, specifically examining such aspects as the core-cusp problem or the baryonic Tully-Fisher relation. For a wide range of parameters, we expect the small-scale phenomenology of the condensate to be similar to the superfluid DM case, thus avoiding usual astrophysical constraints on fermionic particle dark matter candidates [74]. Other phenomena associated with superfluid DM, such as vortices [90, 91, 92], are also possible and worth exploring in this context. It will be important to distinguish this model from a fundamental ultralight scalar dark matter candidate. One possibility for future work is to study the possible neutrino oscillations induced by the background neutrino condensate as neutrinos propagate through it. From a more theoretical perspective, one could also investigate how the four-fermion interaction affects the fermionic densities and get a fully-fledged expression for the condensate contribution to the dark matter density. Another direction would be explicitly constructing the baryogenesis model that gives origin to a degenerate fermion background with suitable chemical potential.
@@ Appendix A Fierz rearrangement
$\Gamma^{A}$ =$=\{\mathbb{1},\gamma^{\mu},\gamma^{\mu\nu},\gamma^{\mu\nu\rho},\gamma^{\mu\nu\rho\sigma}\},$
$\Gamma_{A}$ =$=\{\mathbb{1},\gamma_{\mu},\gamma_{\nu\mu},\gamma_{\rho\nu\mu},\gamma_{\sigma\rho\nu\mu}\},\quad\sigma>\rho>\nu>\mu,$ (42)

### v3

@@ V Discussion and Conclusion
In this work we have presented an analysis regarding whether dark matter states are viable within the structure of the Standard Model. Although it is not possible, due to precision data, to construct dark states using only Standard Model fields, it is possible to do so within the gauge structure of the Standard Model, by augmenting the matter sector with well-motivated additional fermionic fields, e.g. the sterile neutrino in this work. Other scenarios would have their own unique aspects, but much of the physics common to this and other scenarios arises from the well-known physics of superconductivity.
@@
The necessary chemical potential in the neutrino case is adjusted to account for the fact that the neutrino mass is bounded from above and that the interaction is mediated by the Higgs which is relatively heavy for such an interaction. Due to the constraints on the neutrino mass and the Higgs mass, the possible values for the DM particle are fixed primarily by the chemical potential and it is thus possible in this case to generate DM masses in the ultralight range.
@@
An important aspect left for future work is the phenomenology of this model, specifically examining such aspects as the core-cusp problem or the baryonic Tully-Fisher relation. For a wide range of parameters, we expect the small-scale phenomenology of the condensate to be similar to the superfluid DM case, thus avoiding usual astrophysical constraints on fermionic particle dark matter candidates [42]. Other phenomena associated with superfluid DM, such as vortices [59, 60, 61], are also possible and worth exploring in this context. It will be important to distinguish this model from a fundamental ultralight scalar dark matter candidate. One possibility for future work is to study the possible neutrino oscillations induced by the background neutrino condensate as neutrinos propagate through it. Another possibility for implementing the condensate from attractive channels within the SM arises for the case of the color-flavor locking phase of VLQs due to single-gluon exchange. We leave the exploration of this possible scenario to a future work. From a more theoretical perspective, one could also investigate how the four-fermion interaction affects the fermionic densities and get a fully-fledged expression for the condensate contribution to the dark matter density. Another direction would be explicitly constructing the baryogenesis model that gives origin to a degenerate fermion background with suitable chemical potential.
@@ Appendix A Fierz rearrangement
$\Gamma^{A}$ $=\{\mathbb{1},\gamma^{\mu},\gamma^{\mu\nu},\gamma^{\mu\nu\rho},\gamma^{\mu\nu\rho\sigma}\},$
$\Gamma_{A}$ $=\{\mathbb{1},\gamma_{\mu},\gamma_{\nu\mu},\gamma_{\rho\nu\mu},\gamma_{\sigma\rho\nu\mu}\},\quad\sigma>\rho>\nu>\mu,$ (34)